OpenStudy (anonymous):

Identify the range of the equation y = −x2 − 8x − 10. All real numbers y ≥ −2 y ≤ 6 y ≤ −4

OpenStudy (anonymous):

\[x ^{2}+8x+10+y=0\] \[for x be real D \ge 0, 8^{2}-4\left( 10+y \right) \ge 0\] \[64-40-4y \ge 0,24-4y \ge 0,-4y \ge -24, y \le 6\] \[Range is y \le 6\]